OpenStudy (anonymous):

The sum of three consecutive odd numbers is 153. Which is the MIDDLE number? options are : A-47 B-49 C-51 D-53

OpenStudy (cwrw238):

Hint : the middle number will be the mean of the 3 numbers How do you find the mean of 3 numbers?

OpenStudy (anonymous):

Mean?

OpenStudy (cj49):

N-2,N,N+2 N-2+N+N+2=153 3N=153 N=51

OpenStudy (cwrw238):

average 153 / 3 = 51 3 numbers are 50,51,52

OpenStudy (cwrw238):

oh! sorry - i misread the question - i took it to mean 3 consecutive numbers - not consecutive odd

OpenStudy (anonymous):

okay i was confused?

OpenStudy (cj49):

the ans is 51

OpenStudy (cwrw238):

- though the answer happens to be the same

OpenStudy (cj49):

49+51+53=153